OpenStudy (anonymous):

Which of the x values are solutions to the inequality 4(2 – x) > –2x – 3(4x + 1)? Check all that apply. x = –1.1 x = –2.2 x = 0 x = –10 x = 10

OpenStudy (sleepyhead314):

Hello @Roderich and Welcome to OpenStudy! :) first use the distributive property to simplify what they gave you for example 4(2 - x) = 4*2 - 4*x = 8 - 4x can you do that? :)

OpenStudy (kittiwitti1):

mero, what the calculator said o vo \[4(2-x)=4\times2 - 4\times x = 8-4x\]